OpenStudy (anonymous):

Prove by induction that for all positive integers, 5^n>= n^2

OpenStudy (anonymous):

so, assume that \[5^k \ge k^2 \] is true, if we can show that the statement is true for k+1, then the statement is true for all k \[5^(k+1)\ge (k+1)^2\] \[5^k * 5 \ge k^2 +2 k +1 \] where do we go from here

ganeshie8 (ganeshie8):

\(5^k \ge k^2\) multiply both sides by 5 \(5^{k+1} \ge 5k^2\) \(~~~~~~~ \ge k^2 + 2k^2 + 2k^2\) \(~~~~~~~ \ge k^2 + 2k + 1\) \(~~~~~~~ \ge (k+1)^2\)

OpenStudy (anonymous):

5^(k+1)=5^k *5= 5^k+4*5^k 4*5^k>4k^2 (as per assumption) > 2k^2+1

OpenStudy (anonymous):

sorry, 4k^2>2k+1

OpenStudy (anonymous):

combine those

ganeshie8 (ganeshie8):

if you don't like the earlier method, you may use a sub proof to show 4k^2 > 2k+1

ganeshie8 (ganeshie8):

show that 4k^2 > 2k+1 before diving into induction step

OpenStudy (anonymous):

that's easy, for integers, 4k^2>=4k>=3k+1>2k+1

OpenStudy (anonymous):

for positive integers*

OpenStudy (anonymous):

thank guys

ganeshie8 (ganeshie8):

yw !

OpenStudy (anonymous):

So do we assume \(k>0\) then?

OpenStudy (anonymous):

Since it is all positive integers, then \(n\geq 1\). So I guess it checks out.